OpenStudy (anonymous):

The base is 15cm. The height is 13cm. The distance from the right altitude to the vertex is 7cm.

OpenStudy (anonymous):

 Find the area of the figure. Round your answer to the nearest tenth.

OpenStudy (anonymous):

A = 1/2 bh h = 13 b = 15 + 7 A = 1/2 (13)(22) A = 143cm

OpenStudy (anonymous):

lol you still take this?

OpenStudy (anonymous):

yea:/

OpenStudy (anonymous):

Find the area of the whole shape! Idk how to do this crap.

OpenStudy (anonymous):

just search for the formula and plug in

OpenStudy (anonymous):

that easy